Express each angle measure to radian measure.
a.)5 degrees 5 minutes and 5 seconds
b.)3\5 rev.
c.)1 degree 2 minutes 3 seconds
d.)4.1 rev.
e.)60 degrees 5 minutes 28 seconds
Express each angle measure to degrees measure.
a.)1.2 rad
b.)5\3 rev.
c.)8 rad
d.)1.4 rev
e.)4.2 rad

Answer by stanbon(75887)

Express each angle measure to radian measure.
a.)5 degrees 5 minutes and 5 seconds
= 5+5/60+5/3600 = 5.0847222222 degrees
= 5.0847222222 * (pi/180) = 0.0887451443... radians
------------------
b.)3\5 rev.* 2(pi) = 3.7699... radians
---------------------
c.)1 degree 2 minutes 3 seconds
Use same pattern as "a"
------------------------
d.)4.1 rev.
Use same pattern as "b"
-----------------------
e.)60 degrees 5 minutes 28 seconds
Use same pattern as "a"
---------------------
Express each angle measure to degrees measure.
a.)1.2 rad *(180/pi) = 68.7549... degrees
b.)5\3 rev.* 360 = 600 degrees
c.)8 rad
Use same pattern as "a"
d.)1.4 rev
Use same pattern as "b"
e.)4.2 rad
Use same pattern as "a"
==========================
Cheers,
Stan H.